OpenStudy (anonymous):

Integration using tables help please. Integrate sqrt(9-25x^2)dx

OpenStudy (anonymous):

So I am using the formula int sqrt(a^2-x^2)dx = x/2*sqrt(a^2-x^2)+a^2/2sin^-1 x/a +c where a=3 and u=5x

OpenStudy (anonymous):

For an integral of the form \[\int\sqrt{a-bx^2}~dx\] you'd use a substitution, \(x=\sqrt{\dfrac{a}{b}}\sin u\), then \(dx=\sqrt{\dfrac{a}{b}}\cos u~du\). \[\int\sqrt{a-b\left(\sqrt{\dfrac{a}{b}}\sin u\right)^2}\left(\sqrt{\dfrac{a}{b}}\cos u~du\right)\] Simplifying, you have \[\sqrt{\frac{a}{b}}\int\sqrt{a-b\left(\frac{a}{b}\sin^2 u\right)}\cos u~du\\ \sqrt{\frac{a}{b}}\int\sqrt{a-a\sin^2 u}\cos u~du\\ \sqrt{\frac{a}{b}}\sqrt a\int\sqrt{1-\sin^2 u}\cos u~du\\ \frac{a}{\sqrt b}\int\sqrt{\cos^2 u}\cos u~du\\ \frac{a}{\sqrt b}\int\cos^2 u~du\] Half-angle identity: \[\cos^2u=\frac{1}{2}\left(1+\cos2u\right)\] So, you get \[\frac{a}{2\sqrt b}\int(1+\cos2u)~du\\ \frac{a}{2\sqrt b}\left(u+\frac{1}{2}\sin2u\right)+C\] Now you have to back-substitute: \[x=\sqrt{\frac{a}{b}}\sin u~~\iff~~\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)=u\] Let's also rewrite the \(\sin 2u\) using the double angle identity: \[\sin2u=2\sin u\cos u\] So, we have \[\frac{a}{2\sqrt b}\left(u+\sin u\cos u\right)+C\\ \frac{a}{2\sqrt b}\left[\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)+\sin \left(\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)\right)\cos \left(\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)\right)\right]+C\\ \frac{a}{2\sqrt b}\left[\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)+\sqrt{\frac{b}{a}}x\cos \left(\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)\right)\right]+C\] The cosine term is a bit tricky, but you can draw a reference triangle to determine the ratio:  which gives you \[\cos u=\cos\left(\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)\right)=\sqrt\frac{a-bx^2}{a}\] So, you get \[\frac{a}{2\sqrt b}\left[\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)+\sqrt{\frac{b}{a}}\sqrt\frac{a-bx^2}{a}~x\right]+C\\ \frac{a}{2\sqrt b}\left[\sin^{-1}\left(\sqrt{\frac{b}{a}}x\right)+\frac{\sqrt b}{a}\sqrt{a-bx^2}~x\right]+C\] So with \(a=9\) and \(b=25\), \[\begin{align*}\int\sqrt{9-25x^2}~dx&=\frac{9}{2\sqrt{25}}\left[\sin^{-1}\left(\sqrt{\frac{25}{9}}~x\right)+\frac{\sqrt {25}}{9}\sqrt{9-25x^2}~x\right]+C\\ &=\frac{9}{10}\left[\sin^{-1}\left(\frac{5}{3}x\right)+\frac{5}{9}\sqrt{9-25x^2}~x\right]+C\\ &=\frac{9}{10}\sin^{-1}\left(\frac{5}{3}x\right)+\frac{x}{2}\sqrt{9-25x^2}+C \end{align*}\]